OpenStudy (barrelracing):

Steve manages a skate park and he has been analyzing the attendance data. Steve finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 5x. Steve also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by f(x) = −x + 6. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents.

OpenStudy (anonymous):

Hi

OpenStudy (baru):

graph?

OpenStudy (barrelracing):

sorry here it is @baru

OpenStudy (anonymous):

Hi

OpenStudy (barrelracing):

hi

OpenStudy (anonymous):

Hru

OpenStudy (barrelracing):

just trying to get this done can you help?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

I'm 18

OpenStudy (anonymous):

Let me read the questions

OpenStudy (anonymous):

Omg

OpenStudy (anonymous):

Anyone

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

This questions hard

OpenStudy (baru):

x: temperature f(x): number of people who are in the park g(x): number of people leaving early. if we solve them, we set f(x)=g(x) this means, the point of intersection is the temperature at which ALL the people in the park leave early

OpenStudy (anonymous):

Oo

OpenStudy (anonymous):

I'm not school today

OpenStudy (barrelracing):

so the answer would be f(x)=g(x)?

OpenStudy (anonymous):

I going school on tomorrow

OpenStudy (baru):

you need to solve f(x)=g(x) i.e 5^x = -x+6

OpenStudy (anonymous):

That easy

OpenStudy (barrelracing):

ok hold on

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

@baru hi

OpenStudy (baru):

to solve them, just look at the graph, see where they intersect

OpenStudy (anonymous):

Omg

OpenStudy (anonymous):

If I was on my computer

OpenStudy (anonymous):

I could slove this 😥😥@baru

OpenStudy (barrelracing):

(1,5)

OpenStudy (barrelracing):

@baru

OpenStudy (anonymous):

He offline

OpenStudy (baru):

yes, thats right

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Hi